# Dispersion of first sound in a weakly interacting ultracold Fermi liquid: an exact calculation

### v2

@@ Abstract
At low temperature, a normal gas of unpaired spin-1/2 fermions is one of the cleanest realizations of a Fermi liquid. It is described by Landau’s theory, where no phenomenological parameters are needed as the quasiparticle interaction function can be computed perturbatively in powers of the scattering length $a$, the sole parameter of the short-range interparticle interactions. Obtaining an accurate solution of the transport equation nevertheless requires a careful treatment of the collision kernel, as the uncontrolled error made by the relaxation time approximations increases when the temperature $T$ drops below the Fermi temperature. Here, we study sound waves in the hydrodynamic regime up to second order in the Chapman-Enskog’s expansion. We find that the frequency $\omega_{q}$ of the sound wave is shifted above its linear depart as $\omega_{q}=c_{1}q(1+\alpha q^{2}\tau^{2})$ where $c_{1}$ and $q$ are the speed and wavenumber of the wave and the typical collision time $\tau$ scales as $1/a^{2}T^{2}$. Besides the shear viscosity, the coefficient $\alpha$ is described by a single second-order collision time which we compute exactly from an analytical solution of the transport equation, resulting in a positive dispersion $\alpha>0$. Our results suggest that ultracold atomic Fermi gases are an ideal experimental system for quantitative tests of second order hydrodynamics.
@@
Introduction— Landau’s Fermi liquid theory is an effective theory which, when it is applicable, greatly simplifies the description of the many-body dynamics of a system of fermions, breaking it down to a single kinetic equation on a distribution of dilute quasiparticles. It is a very successful theory in describing the phenomenology of a wide class of fermionic systems, such as liquid 3He [1, 2], electron gases [3, 4, 5], quantum gases [6], down to nuclear/neutron matter [7, 44]. Nevertheless, if both the the quasiparticle interaction function and the collision probability [8] are known, either from measurements [9, 10] or from a microscopic calculation [11], Fermi liquid theory provides quantitative predictions on dynamical properties, such as the transport coefficients [12, 13, 14].
@@
In normal 3He, despite decades of research, there is still a discrepancy between theory and experiment on the value of the transport coefficients. Consider for instance the shear viscosity $\eta$, which scales with temperature as $T^{-2}$: theory still underestimate the product $\eta T^{2}$ from the measurement by more than $20\%$ [15, 1, 16, 17, 18]. This is due to a limited knowledge of the quasiparticle interaction function and collision probability [1], which were not computed from a microscopic theory, and whose experimental determination is limited to the lowest spherical harmonics. As a consequence, the exact solutions of the transport equation [12, 13, 14] were never validated experimentally, and relaxation time approximations [19, 1] remain in use today [17].
@@
Ultracold gases of fermionic gases provide exciting opportunities to quantitatively test those transport calculations [20]. These gases behave as Fermi liquids when the $s$-wave scattering length $a$ is negative and sufficiently small to open a regime of temperatures $T_{c}\ll T\ll T_{\textrm{F}}$ where $T_{c}$ is the superfluid critical temperature and $T_{\textrm{F}}$ the Fermi temperature. A Fermi liquid regime may also exist when $a$ is positive and small enough to suppress three-body recombination [21]. The quasiparticle dispersion and interaction function can be computed perturbatively in powers of $k_{\textrm{F}}|a|$ (with $k_{\textrm{F}}$ the Fermi wavenumber) [11]. Experimentally, both the interaction strength and the temperature can be varied such that the typical collision time $\tau$ can be adjusted over several orders of magnitude [22, 23], to explore both collisionless and hydrodynamic regimes [20]. Flat-bottom potentials [24, 25], where sound can be excited at very low wavevector $q$ in homogeneous samples allow to study the propagation and attenuation of sound waves in a very controlled environment [23].
@@
Theoretically, great efforts were devoted to the calculation of the viscosity at strong coupling, in particular in the unitary regime $|a|=+\infty$ [26, 27, 28] and exact results are available in the high temperature virial regime [29, 30, 31]. At intermediate temperatures however, a controlled approach has not been found due to the absence of a separation of timescales between the collisional and kinetic dynamics [32]. The temperature range of the Fermi liquid regime shrinks as the quasiparticle cross section increases with the interaction strength [33], and it is eventually hidden by the onset of a superfluid phase at a critical temperature $T_{c}\approx 0.17T_{\textrm{F}}$ at unitarity. Below $T_{c}$, sound attenuation is dominated by phonon-phonon interactions [26, 28, 34]. In the weakly interacting normal phase, the transport coefficients were computed using relaxation time approximations [6, 35, 36, 37], (either in the “variational” [35, 37], or in the original Abrikosov-Khalatnikov [6, 36] formulation), even though their uncontrolled error increases towards low temperatures. In this Letter, we perform an exact calculation of the transport coefficients [12, 13] to lowest order in $k_{\textrm{F}}|a|$, and show that the error of the relaxation time approximations is significant, up to 25%.
@@
For negative values of $a$, the quasiparticle interactions are attractive [38], which prevents the emergence of a zero sound mode as in liquid 3He. We thus lack a parameter similar to $c_{0}-c_{1}$, where $c_{0}$, $c_{1}$ are the speed of zero and first sound respectively, to characterize the dispersion of sound, as was done in 3He [15]. In this work, we derive the leading-order deviation of the frequency $\omega_{q}$ from its linear depart $c_{1}q$. To do so, we solve the transport equation to the order $\tau^{2}$ of the Chapman-Enskog’s expansion. This is the so-called second hydrodynamic order often used in a relativistic context to cure the acausality of the diffusion equations characteristic of dissipative hydrodynamics [39]. We find the exact solution of the transport equation by decomposing the quasiparticle distribution function on a basis of orthogonal polynomials adapted to the low temperature limit [40]. Remarkably, the frequency shift involves only two parameters of the collision kernel: the viscous relaxation time $\tau_{\eta}$ and a second-order viscous time $t_{\eta}$, which we both compute exactly. As for the dispersion of the sound branch [34], we find that $\omega_{q}$ is above its linear depart $c_{1}q$, the deviation being proportional to $q^{3}\tau^{2}$ with $\tau$ scaling as $1/a^{2}T^{2}$.
@@
Transport equation— Landau’s theory postulates that a Fermi liquid is described by a local quasiparticle distribution $n_{\sigma}(\textbf{p},\textbf{r},t)$, which is the number of quasiparticles of spin $\sigma$ having momentum p at position r and time $t$. This distribution slightly deviates from its thermal equilibrium value $n^{\rm eq}_{\sigma}$ due to some slowly-varying (in both space and time) perturbation. The energy of an arbitrary quasiparticle configuration is expanded to second ordrer:
$E=E_{0}+\sum_{\textbf{p},\textbf{r},\sigma}\epsilon_{\rm eq,\sigma}(\textbf{p})\delta n_{\sigma}(\textbf{p},\textbf{r})\\ +\frac{1}{2}\sum_{\textbf{p},\sigma,\textbf{p}^{\prime},\sigma^{\prime},\textbf{r}}f_{\sigma\sigma^{\prime}}(\textbf{p},\textbf{p}^{\prime})\delta n_{\sigma}(\textbf{p},\textbf{r})\delta n_{\sigma^{\prime}}(\textbf{p}^{\prime},\textbf{r})$ (1)
@@
where $\delta n_{\sigma}=n_{\sigma}-n^{\rm eq}_{\sigma}$ is the fluctuation about thermal equilibrium $n_{\sigma}^{\rm eq}(\epsilon_{\rm eq})=1/(1+\textrm{e}^{(\epsilon_{\rm eq}-\mu)/T})$. In the general case, the equilibrium dispersion relation $\epsilon_{\rm eq}$ and interaction function $f_{\sigma\sigma^{\prime}}$ are phenomenological parameters usually rexpressed in terms of an effective mass and Landau parameters. In the case of a weakly interacting gas with contact interactions, these quantities can be calculated perturbatively in powers of the coupling constant $g=4\pi a/m$ [11] (we use $\hbar=k_{\rm B}=1$ throughout this work). To first order in perturbation theory, we have:
@@
$\epsilon_{\sigma}^{\rm eq}(\textbf{p})=\frac{p^{2}}{2m}+g\rho_{{\rm eq},\sigma^{\prime}},\quad f_{\uparrow\downarrow}={g}/V,\quad f_{\sigma\sigma}=0$ (2)
where $V$ is the volume of the gas.
The evolution of the quasiparticle distribution is described by a transport equation:
$\frac{\partial n_{\sigma}}{\partial t}+\frac{\partial\epsilon_{\sigma}}{\partial\textbf{p}}\cdot\frac{\partial n_{\sigma}}{\partial\textbf{r}}-\frac{\partial(\epsilon_{\sigma}+U_{\sigma})}{\partial\textbf{r}}\cdot\frac{\partial n_{\sigma}}{\partial\textbf{p}}=I_{\rm coll,\sigma}$ (3)
@@
where $U_{\sigma}$ is an external driving field, $\epsilon_{\sigma}(\textbf{p},\textbf{r})=\epsilon_{\sigma}^{\rm eq}(\textbf{p})+\sum_{\textbf{p}^{\prime},\sigma^{\prime}}f_{\sigma\sigma^{\prime}}(\textbf{p},\textbf{p}^{\prime})\delta n_{\sigma^{\prime}}(\textbf{p}^{\prime},\textbf{r})$ is the local energy of the quasiparticles, and the collision integral is given in this weakly interacting limit by Fermi’s golden rule
@@
As we seek the eigenmodes of the transport equation, we assume that the drive is weak and linearize Eq. (3) around equilibrium. We focus here on the unpolarized case so $\mu$ is the common chemical potential of the two spin species. We also restrict ourselves to excitations of the total density, and define the total driving field $U_{\rm tot}=U_{\uparrow}+U_{\downarrow}$, that we decompose in Fourier space $U_{\rm tot}(\textbf{r},t)=\text{Re}(U\sum_{\textbf{q}}\textrm{e}^{\textrm{i}(\textbf{q}\cdot\textbf{r}-\omega t)})$. In Fourier space, the transport equation obeyed by $\delta n=\delta n_{\uparrow}+\delta n_{\downarrow}$ becomes:
$\left(\omega-\frac{\textbf{p}\cdot\textbf{q}}{m}\right)\delta n(\textbf{p})+\frac{\partial n_{\rm eq}}{\partial\epsilon_{\rm eq}}\frac{\textbf{p}\cdot\textbf{q}}{m}\left(g\delta\rho+U\right)=\textrm{i}I_{\rm lin}$ (5)
where $\delta\rho(\textbf{q},\omega)=(1/V)\sum_{\textbf{p}^{\prime}}\delta n(\textbf{p}^{\prime},\textbf{q},\omega)$ is the fluctuation of the total density about $\rho_{\rm eq}=\rho_{\rm eq,\uparrow}+\rho_{\rm eq,\downarrow}$, and the linearized collision integral $I_{\rm lin}(\{\delta n(\textbf{p}^{\prime})\}_{\textbf{p}^{\prime}})=I_{\rm lin,\uparrow}+I_{\rm lin,\downarrow}$ is obtained by linearizing Eq. (4) with respect to the function $n_{\sigma}(\textbf{p})$.
@@
$\delta n(\textbf{p})=U\frac{\partial n_{\rm eq}}{\partial\epsilon_{\rm eq}}\nu(\epsilon,\theta)\text{ with }\epsilon=\frac{\epsilon_{\rm eq}-\mu}{T}=\frac{p^{2}/2m-\epsilon_{\textrm{F}}}{T}$ (6)
where we have used the equation of state $\mu=\epsilon_{\textrm{F}}+g\rho_{\rm eq}/2$ of the weakly interacting Fermi gas. Note that the reduced energy $\epsilon$ varies from $\infty$- ∞ to +$\infty$+ ∞ as the momentum $p$ varies from 00 to +$\infty$+ ∞.
Restricting the transport equation to momenta $p$ lying in the relevant energy shell around the Fermi surface, we obtain, to leading order in $T/T_{\textrm{F}}$:
$(c-\cos\theta)\nu(\epsilon,\theta)-\frac{k_{\textrm{F}}a}{\pi}\cos\theta\int_{-\infty}^{+\infty}\textrm{d}\epsilon^{\prime}g(\epsilon^{\prime})\int_{0}^{\pi}\sin\theta^{\prime}\textrm{d}\theta^{\prime}\nu(\epsilon^{\prime},\theta^{\prime})\\ +\frac{\textrm{i}}{\pi\omega_{0}\tau}\left[\Gamma(\epsilon)\nu(\epsilon,\theta)+\int_{-\infty}^{+\infty}\textrm{d}\epsilon^{\prime}\int\sin\theta^{\prime}\textrm{d}\theta^{\prime}\frac{\textrm{d}\phi^{\prime}}{2\pi}\mathcal{N}_{\rm od}(\epsilon,\epsilon^{\prime},u)\nu(\epsilon^{\prime},\theta^{\prime})\right]=-\cos\theta$ (7)
@@
$\tau=\frac{1}{2ma^{2}T^{2}}$ (8)
is a typical collision time. The diagonal part of the collision kernel is given by
@@
This dimensionless function sets the physical quasiparticle lifetime to $\tau_{\rm qp}(\epsilon)=\pi\tau/\Gamma(\epsilon)$. The off-diagonal part of the kernel reads
@@
The hydrodynamic limit— The hydrodynamic limit is the regime of short collision times

### v3

@@ Abstract
At low temperature, a normal gas of unpaired spin-1/2 fermions is one of the cleanest realizations of a Fermi liquid. It is described by Landau’s theory, where no phenomenological parameters are needed as the quasiparticle interaction function can be computed perturbatively in powers of the scattering length $a$, the sole parameter of the short-range interparticle interactions. Obtaining an accurate solution of the transport equation nevertheless requires a careful treatment of the collision kernel, as the uncontrolled error made by the relaxation time approximations increases when the temperature $T$ drops below the Fermi temperature. Here, we study sound waves in the hydrodynamic regime up to second order in the Chapman-Enskog’s expansion. We find that the frequency $\omega_{q}$ of the sound wave is shifted above its linear departure as $\omega_{q}=c_{1}q(1+\alpha q^{2}\tau^{2})$ where $c_{1}$ and $q$ are the speed and wavenumber of the sound wave and the typical collision time $\tau$ scales as $1/a^{2}T^{2}$. Besides the shear viscosity, the coefficient $\alpha$ is described by a single second-order collision time which we compute exactly from an analytical solution of the transport equation, resulting in a positive dispersion $\alpha>0$. Our results suggest that ultracold atomic Fermi gases are an ideal experimental system for quantitative tests of second-order hydrodynamics.
I Introduction
@@
Landau’s Fermi liquid theory is an effective theory which, when it is applicable, greatly simplifies the description of the many-body dynamics of a system of fermions, breaking it down to a single kinetic equation on a distribution of dilute quasiparticles. It is a very successful theory in describing the phenomenology of a wide class of fermionic systems, such as liquid 3He [1, 2], electron gases [3, 4, 5], quantum gases [6], down to nuclear/neutron matter [7, 8]. Nevertheless, if both the quasiparticle interaction function and the collision probability [9] are known, either from measurements [10, 11] or from a microscopic calculation [12], Fermi liquid theory provides quantitative predictions on dynamical properties, such as the transport coefficients [13, 14, 15].
@@
In normal 3He, despite decades of research, there is still a discrepancy between theory and experiment on the value of the transport coefficients. Consider for instance the shear viscosity $\eta$, which scales with temperature as $T^{-2}$: the theory still underestimates the product $\eta T^{2}$ from the measurement by more than $20\%$ [16, 1, 17, 18, 19]. This is due to a limited knowledge of the quasiparticle interaction function and collision probability [1], which are not computed from a microscopic theory, and whose experimental determination is limited to the lowest spherical harmonics. As a consequence, the exact solutions of the transport equation [13, 14, 15] were never validated experimentally, and relaxation time approximations [20, 1] remain in use today [18].
@@
Ultracold gases of fermionic gases provide exciting opportunities to quantitatively test those transport calculations [21]. These gases behave as Fermi liquids when the s-wave scattering length $a$ is negative and sufficiently small to open a regime of temperatures $T_{c}\leq T\ll T_{\textrm{F}}$ where $T_{c}$ is the superfluid critical temperature and $T_{\textrm{F}}$ the Fermi temperature. A Fermi liquid regime may also exist when $a$ is positive and small enough to suppress three-body recombination [22]. The quasiparticle dispersion, interaction function and collision probabilities can be computed perturbatively in powers of $k_{\textrm{F}}|a|$ (with $k_{\textrm{F}}$ the Fermi wavenumber) [12]. Experimentally, both the interaction strength and the temperature can be varied such that the typical collision time $\tau$ can be adjusted over several orders of magnitude [23, 24], to explore both collisionless and hydrodynamic regimes [21]. Flat-bottom potentials [25, 26], where sound can be excited at very low wavevector $q$ in homogeneous samples, give access to the propagation and attenuation of sound waves in a very controlled environment [24].
@@
Theoretically, great efforts were devoted to the calculation of the viscosity at strong coupling, in particular in the unitary regime $|a|=+\infty$ [27, 28, 29], and exact results are available in the high-temperature virial regime [30, 31, 32, 33, 34]. At intermediate temperatures however, a controlled approach has not been found due to the absence of a separation of timescales between the collisional and kinetic dynamics [35]. The temperature range of the Fermi liquid regime shrinks as the quasiparticle cross section increases with the interaction strength [36], and it is eventually hidden by the onset of a superfluid phase at a critical temperature $T_{c}\approx 0.17T_{\textrm{F}}$ at unitarity. At temperatures low compared to $T_{c}$, sound attenuation is dominated by phonon-phonon interactions [27, 29, 37].
In the weakly interacting normal phase, the transport coefficients were computed using relaxation time approximations [6, 38, 39, 40] (either in the “variational” [38, 40], or in the original Abrikosov-Khalatnikov [6, 39] formulation), even though their uncontrolled error increases toward low temperatures. In this article, we perform an exact calculation of the transport coefficients [13, 14] to lowest order in $k_{\textrm{F}}|a|$, and show that the error of the relaxation time approximations is significant, up to 25%.
@@
For negative values of $a$, the quasiparticle interactions are attractive [41], which prevents the emergence of a zero sound mode as in liquid 3He. We thus lack a parameter similar to $c_{0}-c_{1}$, where $c_{0}$, $c_{1}$ are the speed of zero and first sound respectively, to characterize the dispersion of sound, as was done in 3He [16]. In this work, we derive the leading-order deviation of the frequency $\omega_{q}$ from its linear departure $c_{1}q$. To do so, we solve the transport equation to the order $\tau^{2}$ of the Chapman-Enskog’s expansion. This is the so-called second-order hydrodynamics [42, 43, 44] often used in a relativistic context to cure the acausality of the diffusion equations characteristic of dissipative hydrodynamics [45]. We find the exact solution of the transport equation by decomposing the 3D quasiparticle distribution function on a basis of orthogonal polynomials adapted to the low-temperature limit [46]. Remarkably, the frequency shift involves only two parameters of the collision kernel: the viscous relaxation time $\tau_{\eta}$ and a second-order viscous time $t_{\eta}$, which we both compute exactly. As for the dispersion of the sound branch [37], we find that $\omega_{q}$ is above its linear departure $c_{1}q$, the deviation being proportional to $q^{3}\tau^{2}$ with $\tau$ scaling as $1/a^{2}T^{2}$.
II Transport equation at low temperature
@@
Landau’s theory postulates that a Fermi liquid is described by a local quasiparticle distribution $n_{\sigma}(\textbf{p},\textbf{r},t)$, which is the number of quasiparticles of spin $\sigma$ having momentum p at position r and time $t$. This distribution deviates slightly on average from its value in the Fermi sea $n^{0}_{\sigma}(p)=\Theta(p_{\rm F}-p)$ (where $p_{\rm F}$ is the Fermi momentum and $\Theta$ the Heaviside distribution), and the energy of an arbitrary quasiparticle configuration is expanded to second ordrer in $\delta n_{\sigma}^{0}=n_{\sigma}-n^{0}_{\sigma}$:
$E=E_{0}+\sum_{\textbf{p},\textbf{r},\sigma}\epsilon_{\sigma}^{0}(\textbf{p})\delta n_{\sigma}^{0}(\textbf{p},\textbf{r})\\ +\frac{1}{2}\sum_{\textbf{p},\sigma,\textbf{p}^{\prime},\sigma^{\prime},\textbf{r}}f_{\sigma\sigma^{\prime}}(\textbf{p},\textbf{p}^{\prime})\delta n_{\sigma}^{0}(\textbf{p},\textbf{r})\delta n_{\sigma^{\prime}}^{0}(\textbf{p}^{\prime},\textbf{r})$ (1)
@@
In the general case, the ground state dispersion relation $\epsilon_{\sigma}^{0}$ of quasiparticles and the interaction function $f_{\sigma\sigma^{\prime}}$ are phenomenological parameters usually reexpressed in terms of an effective mass and Landau parameters. In the case of a weakly interacting gas with contact interactions, these quantities can be calculated perturbatively in powers of the coupling constant $g=4\pi a/m$ [12] (we use $\hbar=k_{\rm B}=1$ throughout this work). To first order in perturbation theory, we have:
@@
$\epsilon_{\sigma}^{\rm 0}(\textbf{p})=\frac{p^{2}}{2m}+g\rho_{\sigma^{\prime}},\quad f_{\uparrow\downarrow}={g}/V,\quad f_{\sigma\sigma}=0$ (2)
where $V$ is the volume of the gas, and $\rho_{\sigma}$ the density of fermions of spin $\sigma$.
The time evolution of the quasiparticle distribution is described by a transport equation:
$\frac{\partial n_{\sigma}}{\partial t}+\frac{\partial\epsilon_{\sigma}}{\partial\textbf{p}}\cdot\frac{\partial n_{\sigma}}{\partial\textbf{r}}-\frac{\partial(\epsilon_{\sigma}+U_{\sigma})}{\partial\textbf{r}}\cdot\frac{\partial n_{\sigma}}{\partial\textbf{p}}=I_{\sigma}$ (3)
@@
where $U_{\sigma}$ is an external driving field, $\epsilon_{\sigma}(\textbf{p},\textbf{r})=\epsilon_{\sigma}^{\rm 0}(\textbf{p})+\sum_{\textbf{p}^{\prime},\sigma^{\prime}}f_{\sigma\sigma^{\prime}}(\textbf{p},\textbf{p}^{\prime})\delta n_{\sigma^{\prime}}^{0}(\textbf{p}^{\prime},\textbf{r})$ is the local energy of the quasiparticles, and the collision integral is given in this weakly interacting limit by Fermi’s golden rule
@@
As we seek the eigenmodes of the transport equation, we assume that the drive is weak, slowly varying in both time and space, and we linearize Eq. (3) around the thermal equilibrium distribution $n_{\rm eq}(\epsilon_{0})=1/(1+\textrm{e}^{(\epsilon_{0}-\mu)/T})$. We focus here on the unpolarized case so $\mu$ is the common chemical potential of the two spin species and $\epsilon_{0}\equiv\epsilon^{0}_{\uparrow}=\epsilon^{0}_{\downarrow}$. We also restrict ourselves to excitations of the total density, and define the total driving field $U_{\rm tot}=U_{\uparrow}+U_{\downarrow}$, that we decompose in Fourier space $U_{\rm tot}(\textbf{r},t)=\text{Re}(U\sum_{\textbf{q}}\textrm{e}^{\textrm{i}(\textbf{q}\cdot\textbf{r}-\omega t)})$. Restricting to terms of first order in temperature $T$ and the drive intensity $U$, the transport equation obeyed by $\delta n(\textbf{p})=n_{\uparrow}(\textbf{p})+n_{\downarrow}(\textbf{p})-2n_{\rm eq}(\epsilon_{0}(\textbf{p}))$ is, in Fourier space:
$\left(\omega-\frac{\textbf{p}\cdot\textbf{q}}{m}\right)\delta n(\textbf{p},\textbf{q},\omega)+\frac{\partial n_{\rm eq}}{\partial\epsilon_{0}}\frac{\textbf{p}\cdot\textbf{q}}{m}\left(g\delta\rho(\textbf{q},\omega)+U\right)\\ =\textrm{i}I_{\rm lin}$ (5)
where $\delta\rho(\textbf{q},\omega)$ is the fluctuation of the total density about $\rho_{\rm eq}=\rho_{\uparrow}+\rho_{\downarrow}$ (see Eq. (LABEL:deltarho) below), and the linearized collision integral $I_{\rm lin}$ is given by Eq. (44) in Appendix.
@@
$\delta n(\textbf{p})=U\frac{\partial n_{\rm eq}}{\partial\epsilon_{0}}\Big{|}_{\epsilon_{0}=\epsilon_{\rm 0}(\textbf{p})}\nu(\epsilon,\theta)\\ \text{ with }\epsilon=\frac{\epsilon_{\rm 0}(\textbf{p})-\mu}{T}=\frac{p^{2}/2m-\epsilon_{\textrm{F}}}{T}$ (6)
where $\theta$ is the angle between p and q, and we have used the equation of state $\mu=\epsilon_{\textrm{F}}+g\rho_{\rm eq}/2$ of the weakly interacting Fermi gas. At low temperature, the reduced energy $\epsilon$ varies from -$\infty$ to $\infty$ as the momentum $p$ varies from 0 to $\infty$.
Restricting the transport equation to momenta $p$ lying in the relevant energy shell around the Fermi surface, we obtain, in the limit $T/T_{\textrm{F}}\to 0$ at fixed $\epsilon$ and $\epsilon^{\prime}$:
$(c-\cos\theta)\nu(\epsilon,\theta)-\frac{k_{\textrm{F}}a}{\pi}\cos\theta\int_{-\infty}^{+\infty}\textrm{d}\epsilon^{\prime}g(\epsilon^{\prime})\int_{0}^{\pi}\sin\theta^{\prime}\textrm{d}\theta^{\prime}\nu(\epsilon^{\prime},\theta^{\prime})\\ +\frac{\textrm{i}}{\omega_{0}\tau}\left[\Gamma(\epsilon)\nu(\epsilon,\theta)+\int_{-\infty}^{+\infty}\textrm{d}\epsilon^{\prime}\int\sin\theta^{\prime}\textrm{d}\theta^{\prime}\frac{\textrm{d}\phi^{\prime}}{2\pi}\mathcal{N}_{\rm od}(\epsilon,\epsilon^{\prime},u)\nu(\epsilon^{\prime},\theta^{\prime})\right]=-\cos\theta$ (7)
@@
$\tau=\frac{\pi}{2ma^{2}T^{2}}$ (8)
is a typical collision time. For numerical applications, we replace $\tau$ by:
$\tau_{\sigma}=\frac{\pi}{2m\sigma T^{2}}=\frac{\tau}{4\pi}$ (9)
where $\sigma=4\pi a^{2}$ is the scattering cross section. This is because the center of the hydrodynamic to collisionless crossover occurs for $\omega_{0}\tau_{\sigma}\simeq 1$ rather than $\omega_{0}\tau\simeq 1$. The diagonal part of the collision kernel in Eq. (7) is given by
@@
This dimensionless function sets the physical quasiparticle lifetime to $\tau_{\rm qp}(\epsilon)=\tau/\Gamma(\epsilon)$. The off-diagonal part of the kernel reads
@@
III Hydrodynamic limit
The hydrodynamic limit is the regime of short collision times